OpenStudy (anonymous):

5-3x+7x+3=0 show me the answer and the steps

OpenStudy (anonymous):

5-3x+7x+3=0 add 5+3, and -3x+7x and we get 4x+8=0 subtract 8 from both sides 4x=8 divide by 4 on both sides. since we need to get x alone on one side of the equation. and 8/4=2 so x=2

OpenStudy (anonymous):

wrong...............

OpenStudy (anonymous):

i just need the steps for the right answer

OpenStudy (anonymous):

x=-2 he did it write but just left out the minus sighn when subtracting 8

OpenStudy (anonymous):

Possible intermediate steps: 4 x+8 = 0 Subtract 8 from both sides: 4 x = -8 Divide both sides by 4: x = -2

OpenStudy (anonymous):

Right*

OpenStudy (anonymous):

x=-2 above shown work. that is the right answer.\

OpenStudy (anonymous):

K THXS